calculate the root mean square of bromine molecules at 23 degree celcius and 1.00atm.What is the root mean square of the same gas at 23 degree celcius and 1.5atm?
1
Expert's answer
2019-04-29T03:15:01-0400

Calculate the root mean square of bromine molecules at 23 degree celcius and 1.00atm.What is the root mean square of the same gas at 23 degree celcius and 1.5atm?

Solution:

Urms="\sqrt{\smash[b]{3p/ρ}}" ;

1."\rho" =p*M/R*T; M(Br2)=2*80=160 grams/mol; R=0,08206 atm*L/mol*K; T=23+273=296 K;

"\rho"=1*160/0,08206*296=6,587 grams/L;

Urms="\sqrt{\smash[b]{3*1/6,587}}" =0,4554 m/s.

2."\rho"=1,5*160/0,08206*296=9,88 grams/L;

Urms="\sqrt{\smash[b]{3*1,5/9,88}}" =0,4555 m/s.

Answer: 0,4554 m/s; 0,4555 m/s.